what is the LCM of x^2+6x, and x^2-6x
Answer by nerdybill(7384)

what is the LCM of x^2+6x, and x^2-6x
.
factor each:
x^2+6x = x(x+6)
.
x^2-6x = x(x-6)
.
LCM:
x(x+6)(x-6) (this is the answer)
or, you could expand it:
x(x^2-36)
x^3-36x